To determine the kinetic characteristics of an enzyme you used 1 nmol/L of enzyme in a series of assays where you measured the rate of reactions as you varied the concentration of substrate in each assay (Table A). Estimate from a Michaelis-Menten plot approximate values for Vmax, KM, Kcat, and the specificity constant for this enzyme and substrate. (The only information that is given is this paragraph and the table below).

Table 1:

[S] (μM) v (μmol/L/min)
0 0
5 20
10 50
20 75
40 95
80 125
120 135
160 143
200 150

Estimated: Vmax __________

KM __________ Kcat __________

Spec const __________

Graph table 1, then answer the questions above.
Answer #1

The michaelis-menten equation is

Vmax 160 140 120 100 80 60 40 20 Km 0 -20 0 100 150 200 S] 50 O

Vmax is the maxiumum rate.

Km is the substrate concentration at half of Vmax.

Estimated Vmax = 160 \mu mol/L. min.

Km = 25 \mu M.

Given, enzyme concentration ; [E] = 1 nmol/min = 10-3\mumol/min

Now, Kcat = \frac{V_{max}}{[E]} = \frac{160 \: \mu mol/L.min}{10^{-3}\: \mu mol/L}

Or, Kcat = 160 × 103 min-1  = 1.6 ×105 min-1.

Specific constant = \frac{K_{cat}}{K_m}

= \frac{1.6 \times 10^{5}}{25}

= 6.4×103\mumol-1 min-1 .